# Demystifying 5G Polar and LDPC Codes: A Comprehensive Review and Foundations

### v2

@@ II-D Soft-Decision based LDPC decoding Algorithms
$\mathscr{L}\left(\mathrm{r}_{ji}\right)=\left(\prod_{i\prime\in\mathsf{BN}_{j\backslash i}}\operatorname{sign}\left(\mathscr{L}\left(\mathrm{q}_{i^{\prime}j}\right)\right)\right)\cdot\min_{i\prime\in\mathsf{BN}_{j\backslash i}}\left(\left|\mathscr{L}\left(\mathrm{q}_{i^{\prime}j}\right)\right|\right).$ (24)
@@
(1) Update Input LLRs :$\begin{array}[]{l}\texttt{(1) Update Input LLRs }:\\ \quad\mathscr{L}\left(\mathrm{q}_{ij}\right)=\mathscr{L}\left(\mathrm{Q}_{i}\right)-\mathscr{L}\left(\mathrm{r}_{ji}\right).\\ \texttt{(2) Perform CNP}\\ \quad\mathscr{L}^{\prime}\left(\mathrm{r}_{ji}\right)=2\tanh^{-1}\left(\displaystyle\prod_{i^{\prime}\in\mathsf{BN}_{j\backslash i}}\tanh\left(\frac{1}{2}\mathscr{L}\left(\mathrm{q}_{i^{\prime}j}\right)\right)\right).\\ \texttt{(3) Perform Output LLRs}\\ \quad\mathscr{L}^{\prime}\left(\mathrm{Q}_{i}\right)=\mathscr{L}\left(\mathrm{q}_{ij}\right)+\mathscr{L}^{\prime}\left(\mathrm{r}_{ji}\right).\end{array}$ (27)
For each layer, the decoding stage (3) works on the combined input obtained from the current LLR inputs $=\mathscr{L}\left(\mathrm{q}_{ij}\right)$ and the previous layer updates $\mathscr{L}^{\prime}\left(\mathrm{r}_{ji}\right)$.
@@ III-A State-of-art Polar codes
[…]
@@ III-B Foundations and Fundamentals
where $\mathbf{B}_{\mathsf{N}}$ is the bit-reversal permutation matrix, $\mathbf{F}^{\otimes n}$ is the $n-fold$ Kronecker product of $\mathbf{F}$ with itself, defined recursively as
@@
[…]
where $\oplus$ denotes addition in $\mathbb{F}_{2}$. The codeword $[\mathrm{x}_{0},\mathrm{x}_{1},\ldots,\mathrm{x}_{7}]$ is transmitted through the physical channels $\mathrm{W}^{8}$, with channel outputs $[\mathrm{y}_{0},\mathrm{y}_{1},\ldots,\mathrm{y}_{7}]$. This structure, comprising coding and channel transmission, creates the channel $\mathrm{W}^{8}:\mathrm{x}_{0},\mathrm{x}_{1},\ldots,\mathrm{x}_{7}\longmapsto\mathrm{y}_{0},\mathrm{y}_{1},\ldots,\mathrm{y}_{7}$.
@@
    The bit sequence input for a given code block to channel coding is denoted by $c(0),c(1),\ldots,c(\mathrm{K}_{r}-1)$ , where $\mathrm{K}_{r}$ is the number of bits in code block number $r$, and each code block is individually encoded. After the encoding process, the resulting coded bit sequence within the $r-th$ code block is denoted by $d_{r}(0),d_{r}(1),\ldots,d_{r}(\mathrm{N}_{r}-1)$ where $\mathrm{N}_{r}=2^{n}$ (code length of the polar code) determined by the following:
@@
      The interleaved vector $\mathbf{c}^{\prime}$ is assigned to the information set along with the PC bits, while the remaining bits in the $\mathrm{N}$-bit vector $\mathbf{u}$ are frozen. Hence, $\mathbf{u}=u(0),u(1),\ldots,u(\mathrm{N}-1)$ is generated according to the clause 5.3.1.2 [6]. Denote $\mathbf{G}_{\mathsf{N}_{r}}=\left(\mathbf{G}_{2}\right)^{\otimes n}$ as the $n-th$ Kronecker power of matrix $\mathbf{G}_{2}$ , where $\mathbf{G}_{2}=\left[\begin{array}[]{ll}1&0\\ 1&1\end{array}\right]$, the output after encoding $\mathbf{d}_{r}=d_{r}(0),d_{r}(1),\ldots,d_{r}(\mathrm{N}_{r}-1)$ is obtained by $\mathbf{d}_{r}=\mathbf{u}_{r}\mathbf{G}_{\mathsf{N}_{r}}$, where encoding is performed in $\mathbb{F}_{2}$.
@@
      The bits input to the sub-block interleaver are the coded bits $d_{r}(0),d_{r}(1),\ldots,d_{r}(\mathrm{N}_{r}-1)$ The coded bits $d_{r}(0),d_{r}(1),\ldots,d(\mathrm{N}_{r}-1)$ are divided into 32 sub-blocks. The bits output from the sub-block interleaver are denoted as $s_{r}(0),s_{r}(1),\ldots,s_{r}(\mathrm{N}_{r}-1)$.
@@
    The input bit sequence for the code block concatenation block are the sequences $f_{r,k}$ , for $r=0,\ldots,C-1$ and $\mathrm{k}=0,\ldots,E_{r}-1$, where $E_{r}$ is the number of rate matched bits for the $r-th$ code block. The output bit sequence from the code block concatenation block is the sequence $g_{\ell}$ for $\ell=0,\ldots,G-1$. Therefore,
@@ III-D Polar Decoding Algorithms
$\mathrm{h}_{i}\left(\mathrm{y}_{0}^{\mathsf{N}-1},\hat{\mathrm{u}}_{0}^{i-1}\right)=\begin{cases}0&\text{ if }\mathscr{L}_{\mathsf{N}}^{(i)}\left(\mathrm{y}_{0}^{\mathsf{N}-1},\hat{\mathrm{u}}_{0}^{i-1}\right)\geq 0,\\ 1&\text{ otherwise .}\end{cases}$ (42)
@@
The SCL decoder was introduced as an extension of the SC decoder. Rather than sequentially computing hard decisions for each bit, it bifurcates into two parallel SC decoders at every decision stage, with each branch maintaining its path metric continuously updated for each path. It’s demonstrated that a list size of 32 is nearly sufficient to reach the ML bound [43].
What is needed is a path metric, computing the likelihood along the entire path of bits. This path metric is established in the following theorem.
Theorem 1.
[73, Theorem 1].
For a path $\ell$ having bits $\hat{\mathrm{u}}_{0}(\ell),\hat{\mathrm{u}}_{1}(\ell),\ldots,\hat{\mathrm{u}}_{i}(\ell)$ and bit number $i\in 0,1,\ldots,\mathrm{N}-1$ , define the path metric as
@@
$\mathscr{L}_{\mathsf{N}}^{(i)}[\ell]=\ln\left(\frac{\mathrm{W}_{\mathsf{N}}^{(i)}\left(y_{0}^{\mathsf{N}-1},\hat{\mathrm{u}}_{0}^{i-1}[\ell]\mid 0\right)}{\mathrm{W}_{\mathsf{N}}^{(i)}\left(\mathrm{y}_{0}^{\mathsf{N}-1},\hat{\mathrm{u}}_{0}^{i-1}[\ell]\mid 1\right)}\right)$ (49)
@@
IV Numerical Results
We are interested in the short block regime, i.e. blocks of 20 to 100 bits, we will carry out simulations using a 100-bit block (TB+CRC), with different coding rates from low to high, QPSK modulation coupled with polar and LDPC coding.
(a) 5G Polar codes.
(b) 5G LDPC codes.
Figure 11: Block Error Rate, 100 bits(TBs+CRC), R={1/4, 1/2, 3/4}, NR LDPC (BP, LBP with Early Stop Criteria (ESC), MaxIter=10) and NR POLAR, (CA-SCL, L=1, 8, 32), QPSK modulations, over ($4\times 1$) SIMO, on Non Line-of-Sight (NLOS) TDL channel, 300ns of delay spread.
It can be noticed in Figures 11a that layered belief propagation (LBP) for LDPC decoding often outperforms standard belief propagation (BP) decoding due to its ability to exploit the structure of the LDPC code more efficiently. Concretely, LDPC codes are often constructed with a layered structure, which means that the parity-check matrix can be organized into layers of variable nodes and check nodes. Hence, LBP takes advantage of this structure by decoding the code layer by layer. This can lead to more efficient message passing compared to BP based on "flood" decoding, which updates messages across all variable nodes and check nodes simultaneously. Moreover, LBP typically has lower computational complexity. By decoding the code layer by layer, the complexity of message passing at each iteration can be reduced, since only messages within a layer need to be exchanged. Due to its improved convergence properties and efficient message passing, layered belief propagation decoding tends to provide better error correction performance compared to standard BP decoding. This is particularly advantageous in scenarios with higher levels of noise or more challenging channel conditions. In general, the layered structure and specialized message passing algorithms for layered belief propagation make it a preferred choice for LDPC decoding in many practical applications, offering improved performance, reduced complexity, and better convergence properties. However, the effectiveness of LBP versus BP decoding can be influenced by factors such as the code rate, the operating signal-to-noise ratio (SNR) region, the complexity of the LDPC code structure, and the computational resources available for decoding. For instance, at lower code rate, performance gaps between LBP and standard BP are more pronounced. Conversely, in terms of convergence time, the merits of decoding paired with the early stop criterion (ESC) can be exploited with the LBP decoding which is amenable to a very interesting trade-off between reliability and latency.
Besides, in setups using polar codes, the CRC-aided successive cancellation list (CA-SCL) decoder is highly regarded for its exceptional error correction capabilities compared to the state of art decoders. The performance of the SCL decoder improves with the number of lists ($L$). Increasing $L$ in the SCL decoder enhances BLER performance, as shown in Figures 11b, but leads to higher decoding complexity, requiring more computational resources.
V Conclusions
In conclusion, this paper has served as a comprehensive guide for practitioners and scholars seeking to understand the channel coding and decoding schemes embedded in the 5G NR standard, with a particular focus on LDPC and polar codes. We began by elucidating the design principles underlying these channel codes, providing essential knowledge from both encoding and decoding perspectives. In parallel, we conducted an extensive literature review to establish the current state of research in this field. Notably, we enriched these foundational discussions with detailed, standard-specific insights that are often difficult to extract from technical specification documents. Through this structured approach, we aim to provide a valuable reference for those interested in the complexities of channel coding within the 5G NR framework.

### v3

@@ 23 $\hat{\mathrm{c}}_{i}=1$ if $\mathscr{L}\left(\mathrm{Q}_{i}\right)<0$ else $\hat{\mathrm{c}}_{i}=0$
[…]
@@
[…]
For each layer, the decoding stage (3) works on the combined input obtained from the current LLR inputs $\mathscr{L}\left(\mathrm{q}_{ij}\right)$ and the previous layer updates $\mathscr{L}^{\prime}\left(\mathrm{r}_{ji}\right)$.
@@ III-A State-of-art Polar codes
Furthermore, the primary polar code decoding algorithms include the SC algorithm [48], the SCL algorithm [73, 74], the CA-SCL algorithm [75], the BP algorithm [76], and the SCAN algorithm [77]. Originally proposed by Arikan, the SC algorithm’s performance diminishes for finite length codes. SCL, an enhancement of SC, offers superior performance by providing multiple paths. CA-SCL, incorporating cyclic redundancy checks on message bits over SCL, significantly boosts performance through simple checksums. Currently, 3GPP polar decoding relies on the CA-SCL algorithm, surpassing LDPC codes. Notably, SC, SCL, and CA-SCL algorithms are hard output, yielding bit sequences rather than LLR values. To facilitate joint designs, soft output algorithms providing LLR values are essential; BP and SCAN algorithms serve this purpose. Decoding delay varies between BP and SCAN: BP utilizes the "flood" rule for message passing [78], while SCAN employs the serial elimination rule for SC-like algorithms. BP exhibits lower decoding delay, whereas SCAN demonstrates superior convergence speed. In terms of performance, these algorithms an be ranked as follows: $\texttt{CA-SCL}>\texttt{state of the art (LDPC, Turbo codes)}>\texttt{SCL}>\texttt{BP}=\texttt{SCAN}>\texttt{SC}$.
@@ III-B Foundations and Fundamentals
where $\mathbf{B}_{\mathsf{N}}$ is the bit-reversal permutation matrix, $\mathbf{F}^{\otimes n}$ is the nfoldn-fold Kronecker product of $\mathbf{F}$ with itself, defined recursively as
@@
$\begin{array}[]{l l }\begin{aligned} &\mathrm{x}_{0}=\mathrm{u}_{0}\oplus\mathrm{u}_{1}\oplus\mathrm{u}_{2}\oplus\mathrm{u}_{4},\\ &\mathrm{x}_{4}=\mathrm{u}_{1}\oplus\mathrm{u}_{3}\oplus\mathrm{u}_{6},\\ &\mathrm{x}_{2}=\mathrm{u}_{2}\oplus\mathrm{u}_{3}\oplus\mathrm{u}_{6},\\ &\mathrm{x}_{6}=\mathrm{u}_{3}\oplus\mathrm{u}_{7},\end{aligned}&\begin{aligned} &\mathrm{x}_{1}=\mathrm{u}_{4}\oplus\mathrm{u}_{5}\oplus\mathrm{u}_{6},\\ &\mathrm{x}_{5}=\mathrm{u}_{5}\oplus\mathrm{u}_{7},\\ &\mathrm{x}_{3}=\mathrm{u}_{6}\oplus\mathrm{u}_{7},\\ &\mathrm{x}_{7}=\mathrm{u}_{7},\end{aligned}\end{array}$ (32)
where $\oplus$ denotes addition in $\mathbb{F}_{2}$. The codeword $\mathrm{x}_{0},\mathrm{x}_{1},\ldots,\mathrm{x}_{7}]$ is transmitted through the physical channels $\mathrm{W}^{8}$, with channel outputs $\mathrm{y}_{0},\mathrm{y}_{1},\ldots,\mathrm{y}_{7}]$. This structure, comprising coding and channel transmission, creates the channel $\mathrm{W}^{8}:\mathrm{x}_{0},\mathrm{x}_{1},\ldots,\mathrm{x}_{7}\longmapsto\mathrm{y}_{0},\mathrm{y}_{1},\ldots,\mathrm{y}_{7}$.
@@
    The bit sequence input for a given code block to channel coding is denoted by $c(0),c(1),\ldots,c(\mathrm{K}_{r}-1)$ , where $\mathrm{K}_{r}$ is the number of bits in code block number $r$, and each code block is individually encoded. After the encoding process, the resulting coded bit sequence within the rthr-th code block is denoted by $d_{r}(0),d_{r}(1),\ldots,d_{r}(\mathrm{N}_{r}-1)$ where $\mathrm{N}_{r}=2^{n}$ (code length of the polar code) determined by the following:
@@
      The interleaved vector $\mathbf{c}^{\prime}$ is assigned to the information set along with the PC bits, while the remaining bits in the $\mathrm{N}$-bit vector $\mathbf{u}$ are frozen. Hence, $\mathbf{u}=u(0),u(1),\ldots,u(\mathrm{N}-1)$ is generated according to the clause 5.3.1.2 [6]. Denote $\mathbf{G}_{\mathsf{N}_{r}}=\left(\mathbf{G}_{2}\right)^{\otimes n}$ as the nthn-th Kronecker power of matrix $\mathbf{G}_{2}$ , where $\mathbf{G}_{2}=\left[\begin{array}[]{ll}1&0\\ 1&1\end{array}\right]$, the output after encoding $\mathbf{d}_{r}=d_{r}(0),d_{r}(1),\ldots,d_{r}(\mathrm{N}_{r}-1)$ is obtained by $\mathbf{d}_{r}=\mathbf{u}_{r}\mathbf{G}_{\mathsf{N}_{r}}$, where encoding is performed in $\mathbb{F}_{2}$.
@@
      The bits input to the sub-block interleaver are the coded bits $d_{r}(0),d_{r}(1),\ldots,d_{r}(\mathrm{N}_{r}-1)$ . The coded bits $d_{r}(0),d_{r}(1),\ldots,d(\mathrm{N}_{r}-1)$ are divided into 32 sub-blocks. The bits output from the sub-block interleaver are denoted as $s_{r}(0),s_{r}(1),\ldots,s_{r}(\mathrm{N}_{r}-1)$.
@@
    The input bit sequence for the code block concatenation block are the sequences $f_{r,k}$ , for $r=0,\ldots,C-1$ and $\mathrm{k}=0,\ldots,E_{r}-1$, where $E_{r}$ is the number of rate matched bits for the rthr-th code block. The output bit sequence from the code block concatenation block is the sequence $g_{\ell}$ for $\ell=0,\ldots,G-1$. Therefore,
@@ III-D Polar Decoding Algorithms
[…]
@@
The SCL decoder was introduced as an extension of the SC decoder. Rather than sequentially computing hard decisions for each bit, it bifurcates into two parallel SC decoders at every decision stage, with each branch maintaining its path metric continuously updated for each path. It’s demonstrated that a list size of 32 is nearly sufficient to reach the maximum likelihood bound [43].
What is needed is a path metric, computing the likelihood along the entire path of bits. This path metric is established in the following theorem [73, Theorem 1], which states: for a path $\ell$ with bits $\hat{\mathrm{u}}_{0}(\ell),\hat{\mathrm{u}}_{1}(\ell),\ldots,\hat{\mathrm{u}}_{i}(\ell)$, and for bit index $i\in{0,1,\ldots,\mathrm{N}-1}$, the path metric is defined as
@@
$\mathscr{L}_{\mathsf{N}}^{(i)}[\ell]=\ln\left(\frac{\mathrm{W}_{\mathsf{N}}^{(i)}\left(\mathrm{y}_{0}^{\mathsf{N}-1},\hat{\mathrm{u}}_{0}^{i-1}[\ell]\mid 0\right)}{\mathrm{W}_{\mathsf{N}}^{(i)}\left(\mathrm{y}_{0}^{\mathsf{N}-1},\hat{\mathrm{u}}_{0}^{i-1}[\ell]\mid 1\right)}\right)$ (50)
@@
IV Conclusions
In conclusion, this work examined the channel coding and decoding schemes defined in the 5G NR standard, with particular emphasis on LDPC and polar codes. We outlined the design principles underlying these codes, offering essential insights from both encoding and decoding perspectives. In parallel, we conducted an extensive review of the literature to capture the current state of research in this domain. Importantly, we complemented these discussions with detailed, standard-specific explanations that are often difficult to extract directly from technical specification documents. Through this structured approach, the study aims to serve as a valuable reference for those exploring the intricate details of channel coding within the beyond 5G NR framework